# Coderivative-Based Newton Methods in Structured Nonconvex and Nonsmooth Optimization

### v2

@@ 1 Introduction
$-\nabla\varphi(x^{k})=\nabla^{2}\varphi(x^{k})d^{k},\quad k\in{\rm I\!N},$ (1.3)
@@
$\bullet$ Sections 5 and 6 are devoted to the globalization of developed algorithms to solve nonconvex structured optimization problems (1.4). In Section 5, we propose and justify a novel line search method, which can be viewed as a generalized proximal gradient line search. To proceed, we exploit the machinery of forward-backward envelopes (FBE) introduced in [45]. The newly proposed line search algorithm doesn’t involved any generalized second-order constructions and ensures the global linear convergence of iterates. Involving the second-order information in Section 6 allows us to design the globalized coderivative-based Newton method (GCNM), which exhibits the the global convergence with local superlinear rate under the assumptions that are less restrictive compered to the previous developments in [26, 27]. In particular, we don’t impose the positive-definiteness of the generalized Hessian on the entire space ${\rm I\!R}^{n}$ (while only at the solution point in some settings) and don’t restrict ourselves to quadratic convex functions $f$ in (1.4). This dramatically extends the scope of applications.
@@ 2 Preliminaries from Variational Analysis
established by Mordukhovich [38, Theorem 3.6] via his limiting coderivative (2.6) and then labeled as the Mordukhovich criterion in Rockafellar and Wets [55, Theorems 7.40 and 7.43]. Broad applications of this result are based on the robustness and full calculus available for (2.6); see [39, 40, 55] for more details and references. A parallel characterization of the (nonrobust) strong metric subregularity property of a closed-graph mapping $F\colon{\rm I\!R}^{n}\rightrightarrows{\rm I\!R}^{m}$ at $(\bar{x},\bar{y})\in\mbox{\rm gph}\,F$ is given by $\mbox{\rm ker}\,DF(\bar{x},\bar{y})=\{0\}$ via the (nonrobust) graphical derivative (2.7) of $F$ at $(\bar{x},\bar{y})$ and is known as the Levy-Rockafellar criterion; see [9, Theorem 4E.1] for more discussions.
@@
^$\widehat{\partial}\varphi(\bar{x}):=\big{\{}v\in{\rm I\!R}^{n}\;\big{|}\;(v,-1)\in\widehat{N}_{{\rm\text{\rm epi}}\,\varphi}\big{(}\bar{x},\varphi(\bar{x})\big{)}\big{\}},$ (2.9)
@@
$\partial^{2}\varphi(\bar{x})(u)=\big{\{}\nabla^{2}\varphi(\bar{x})u\big{\}}\;\mbox{ for all }\;u\in{\rm I\!R}^{n}.$ (2.13)
@@
$\partial^{2}\varphi(\bar{x})(u)=\partial\big{\langle}u,\nabla\varphi\big{\rangle}(\bar{x})\;\mbox{ for all }\;u\in{\rm I\!R}^{n}$ (2.14)
@@
Let us now review some classes of extended-real-valued functions $\varphi\colon{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ broadly used in the paper. The notion of variational $s$-convexity for extended-real-valued functions is a major one we intend to investigate in this paper. For the general case of $s\in{\rm I\!R}$, it was recently introduced in [53], while its earlier versions for $s\geq 0$ appeared in [51] and were later studied and applied in finite-dimensional spaces in [21, 25, 52], and in infinite-dimensional settings in [22, 23, 41]. The case of $s<0$ is related to prox-regularity, a concept originally introduced in [48] and subsequently developed and widely applied in variational analysis and optimization; see [55]. Before introducing the notion of variational $s$-convexity, we recall that a function $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ is $s$-convex for $s\in{\rm I\!R}$ if the quadratically $s$-shifted function $\varphi-\frac{s}{2}\|\cdot\|^{2}$ is convex. In particular, when $s>0$, the function $\varphi$ is referred to as $s$-strongly convex, while for $s<0$, it is labeled as $(-s)$-weakly convex.
@@
where $U_{\varepsilon}:=\{x\in U\mid\varphi(x)<\varphi(\bar{x})+\varepsilon\}$. When $s\leq 0$, $\varphi$ is said to be $(-s)$-level prox-regular at $\bar{x}$ for $\bar{v}$, while for $s=0$ it is variationally convex, and for $s>0$ is $s$-strongly variationally convex at $\bar{x}$ for $\bar{v}$.
The next proposition taken from [53, Theorem 1] lists the characterizations of variationally convex functions via some properties of their limiting subdifferentials.
@@
$\langle v_{1}-v_{2},u_{1}-u_{2}\rangle\geq s\|u_{1}-u_{2}\|^{2}\quad\text{for all }\;(u_{1},v_{1}),(u_{2},v_{2})\in\text{\rm gph}\partial\varphi\cap(U_{\varepsilon}\times V),$ (2.15)
@@
$\varphi(x)\geq\varphi(u)+\langle v,x-u\rangle+\frac{s}{2}\|x-u\|^{2}\quad\text{for all }\;x\in U,\;(u,v)\in\mbox{\rm gph}\,\partial\varphi\cap(U_{\varepsilon}\times V).$ (2.16)
We say that $\varphi$ is subdifferentially continuous at $\bar{x}$ for $\bar{v}\in\partial\varphi(\bar{x})$ if for any $\varepsilon>0$, there is $\delta>0$ with $|\varphi(x)-\varphi(\bar{x})|<\varepsilon$ whenever $(x,v)\in\mbox{\rm gph}\,\partial\varphi\cap(\mathbb{B}_{\delta}(\bar{x})\times\mathbb{B}_{\delta}(\bar{v}))$. Functions that are both prox-regular and subdifferentially continuous are called continuously prox-regular. Note that the continuous prox-regularity is a robust property, i.e., if $\varphi$ is continuously prox-regular at $\bar{x}$ for $\bar{v}$, then it is also continuously prox-regular at any $(x,v)$ sufficiently close to $(\bar{x},\bar{v})$; see [17, Proposition 3.6.]. This is a major class of extended-real-valued functions in second-order variational analysis, being a common roof for particular collections of functions important for applications as, e.g., convex and $C^{1,1}$ functions, amenable functions, etc.; see [55, Chapter 13].
@@
A function $\varphi$ is said to be prox-bounded if there exists $\lambda>0$ such that $e_{\lambda}\varphi(x)>-\infty$ for some $x\in{\rm I\!R}^{n}$. The supremum over all such $\lambda$ is the threshold $\lambda_{\varphi}$ of prox-boundedness for $\varphi$. Prox-boundedness admits various equivalent descriptions and ensures a favorable behavior of Moreau envelopes and proximal mappings. We recall the following results from [55, Theorem 1.25 and Example 10.32].
@@
Let $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ be l.s.c. and prox-bounded with threshold $\lambda_{\varphi}$. Then for all $\lambda\in(0,\lambda_{\varphi})$, we have:
@@
Next we present some properties of Moreau envelopes and proximal mappings taken from [48, Theorem 4.4].
@@
Let $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ be prox-bounded on ${\rm I\!R}^{n}$ and $r$-level prox-regular at $\bar{x}$ for $\bar{v}\in\partial\varphi(\bar{x})$. Then for all $\lambda\in(0,1/r)$ we find a convex neighborhood $U_{\lambda}$ of $\bar{x}+\lambda\bar{v}$ on which the following hold:
@@
(ii) The proximal mapping $\text{\rm Prox}_{\lambda\varphi}$ from (2.18) is single-valued, monotone, and Lipschitz continuous on $U_{\lambda}$ while satisfying the condition ${\rm Prox}_{\lambda\varphi}(\bar{x}+\lambda\bar{v})=\bar{x}$.
@@
$\nabla e_{\lambda}\varphi(x)=\left(\lambda I+(\partial\varphi)^{-1}\right)^{-1}(x)\quad\text{for all}\quad x\in U_{\lambda}.$
The parameter $\lambda$ can be chosen arbitrarily from $(0,\infty)$ with $U_{\lambda}={\rm I\!R}^{n}$ if $\varphi$ is convex.
@@
Let $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ be continuously prox-regular at $\bar{x}\in\mbox{\rm dom}\,\varphi$ for $0\in\partial\varphi(\bar{x})$. The following assertions are equivalent:
(i) $\varphi$ is strongly variationally convex at $\bar{x}$ for $\bar{v}=0$ with modulus $\sigma>0$.
@@
Then $\varphi$ is twice epi-differentiable at $\bar{x}$ for $v$ if for every $w\in{\rm I\!R}^{n}$ and $\tau_{k}\downarrow 0$ there is a sequence $w^{k}\to w$ with
$\frac{\varphi(\bar{x}+\tau_{k}w^{k})-\varphi(\bar{x})-\tau_{k}\langle v,w^{k}\rangle}{\frac{1}{2}\tau_{k}^{2}}\to d^{2}\varphi(\bar{x},v)(w)\;\mbox{ as }\;k\to\infty.$
Twice epi-differentiability has been recognized as an important concept of second-order variational analysis with numerous applications to optimization; see, e.g., [55]. To formulate the related notion of generalized twice differentiability of functions, we proceed by [52] as follows. Given a quadratic matrix $A\in{\rm I\!R}^{n\times n}$, denote the quadratic form associated with the matrix $A$ by
$q_{A}(w):=\langle w,Aw\rangle\;\mbox{ for all }\;w\in{\rm I\!R}^{n}.$ (2.20)
@@
$\Omega_{\varphi}:=\big{\{}(x,v)\in\mbox{\rm gph}\,\partial\varphi\;\big{|}\;\varphi\;\text{ is generalized twice differentiable at }x\;\text{for }v\big{\}}.$ (2.21)
@@
$\text{\rm gph}\;\partial q\subset\text{\rm gph}\;\partial^{2}\varphi(\bar{x},\bar{v})\quad\text{for any }\;q\in\mathrm{quad}\,\varphi(\bar{x}|\bar{v}).$ (2.22)

### v3

# arXiv:2403.04262v3[math.OC] 01 Mar 2026

@@ 1 Introduction
$\nabla\varphi(x^{k})=\nabla^{2}\varphi(x^{k})d^{k},\quad k\in{\rm I\!N},$ (1.3)
@@
$\bullet$ Sections 5 and 6 are devoted to the globalization of developed algorithms to solve nonconvex structured optimization problems (1.4). In Section 5, we propose and justify a novel line search method, which can be viewed as a generalized proximal gradient line search. To proceed, we exploit the machinery of forward-backward envelopes (FBE) introduced in [44]. The newly proposed line search algorithm doesn’t involve any generalized second-order constructions and ensures the global linear convergence of iterates. Involving the second-order information in Section 6 allows us to design the globalized coderivative-based Newton method (GCNM), which exhibits the global convergence with local superlinear rate under the assumptions that are less restrictive compered to the previous developments in [26, 27]. In particular, we don’t impose the positive-definiteness of the generalized Hessian on the entire space ${\rm I\!R}^{n}$ (while only at the solution point in some settings) and don’t restrict ourselves to quadratic convex functions $f$ in (1.4). This dramatically extends the scope of applications.
@@ 2 Preliminaries from Variational Analysis
established by Mordukhovich [37, Theorem 3.6] via his limiting coderivative (2.6) and then labeled as the Mordukhovich criterion in Rockafellar and Wets [54, Theorems 7.40 and 7.43]. Broad applications of this result are based on the robustness and full calculus available for (2.6); see [38, 39, 54] for more details and references. A parallel characterization of the (nonrobust) strong metric subregularity property of a closed-graph mapping $F\colon{\rm I\!R}^{n}\rightrightarrows{\rm I\!R}^{m}$ at $(\bar{x},\bar{y})\in\mbox{\rm gph}\,F$ is given by $\mbox{\rm ker}\,DF(\bar{x},\bar{y})=\{0\}$ via the (nonrobust) graphical derivative (2.7) of $F$ at $(\bar{x},\bar{y})$ and is known as the Levy-Rockafellar criterion; see [9, Theorem 4E.1] for more discussions.
@@
$\widehat{\partial}\varphi(\bar{x}):=\big\{v\in{\rm I\!R}^{n}\;\big|\;(v,-1)\in\widehat{N}_{{\rm\text{\rm epi}}\,\varphi}\big(\bar{x},\varphi(\bar{x})\big)\big\},$
@@
$\partial^{2}\varphi(\bar{x})(u)=\big\{\nabla^{2}\varphi(\bar{x})u\big\}\;\mbox{ for all }\;u\in{\rm I\!R}^{n}.$
@@
$\partial^{2}\varphi(\bar{x})(u)=\partial\big\langle u,\nabla\varphi\big\rangle(\bar{x})\;\mbox{ for all }\;u\in{\rm I\!R}^{n}$
@@
Let us now review some classes of extended-real-valued functions $\varphi\colon{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ broadly used in the paper. The notion of variational $s$-convexity for extended-real-valued functions is a major one we intend to investigate in this paper. For the general case of $s\in{\rm I\!R}$, it was recently introduced in [52], while its earlier versions for $s\geq 0$ appeared in [50] and were later studied and applied in finite-dimensional spaces in [21, 25, 51], and in infinite-dimensional settings in [22, 23, 40]. The case of $s\leq 0$ is related to prox-regularity, a concept originally introduced in [47] and subsequently developed and widely applied in variational analysis and optimization; see [54]. Before introducing the notion of variational $s$-convexity, we recall that a function $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ is $s$-convex for $s\in{\rm I\!R}$ if the quadratically $s$-shifted function $\varphi-\frac{s}{2}\|\cdot\|^{2}$ is convex. In particular, when $s>0$, the function $\varphi$ is referred to as $s$-strongly convex, while for $s<0$, it is labeled as (s)(-s)-weakly convex.
@@
where $U_{\varepsilon}:=\{x\in U\mid\varphi(x)<\varphi(\bar{x})+\varepsilon\}$. When $s\leq 0$, the property is called (s)(-s)-level prox-regularity of $\varphi$ at $\bar{x}$ for $\bar{v}$; we simply say that $\varphi$ is prox-regular at $\bar{x}$ for $\bar{v}$ if this holds for some $s\leq 0$. The function $\varphi$ is variationally convex at $\bar{x}$ for $\bar{v}$ when it is 0-level prox-regular. When $s>0$, $\varphi$ is $s$-strongly variationally convex at $\bar{x}$ for $\bar{v}$, and is said to be strongly variationally convex if this holds for some $s>0$.
The next proposition taken from [52, Theorem 1] lists the characterizations of variationally convex functions via some properties of their limiting subdifferentials.
@@
$\langle v_{1}-v_{2},u_{1}-u_{2}\rangle\geq s\|u_{1}-u_{2}\|^{2}\quad\text{for all }\;(u_{1},v_{1}),(u_{2},v_{2})\in\mbox{\rm gph}\,\partial\varphi\cap(U_{\varepsilon}\times V),$
@@
$\varphi(x)\geq\varphi(u)+\langle v,x-u\rangle+\frac{s}{2}\|x-u\|^{2}\quad\text{for all }\;x\in U,\;(u,v)\in\mbox{\rm gph}\,\partial\varphi\cap(U_{\varepsilon}\times V).$
We say that $\varphi$ is subdifferentially continuous at $\bar{x}$ for $\bar{v}\in\partial\varphi(\bar{x})$ if for any $\varepsilon>0$, there is $\delta>0$ with $|\varphi(x)-\varphi(\bar{x})|<\varepsilon$ whenever $(x,v)\in\mbox{\rm gph}\,\partial\varphi\cap(\mathbb{B}_{\delta}(\bar{x})\times\mathbb{B}_{\delta}(\bar{v}))$. Functions that are both prox-regular and subdifferentially continuous are called continuously prox-regular. When the level of prox-regularity is specified, we refer to it as continuous $r$-level prox-regularity for some $r\geq 0$. Note that the continuous prox-regularity is a robust property, i.e., if $\varphi$ is continuously prox-regular at $\bar{x}$ for $\bar{v}$, then it is also continuously prox-regular at any $(x,v)$ sufficiently close to $(\bar{x},\bar{v})$; see [17, Proposition 3.6.]. This is a major class of extended-real-valued functions in second-order variational analysis, being a common roof for particular collections of functions important for applications as, e.g., convex and $\mathcal{C}^{1,1}$ functions, amenable functions, etc.; see [54, Chapter 13].
@@
A function $\varphi$ is said to be prox-bounded if there exists $\lambda>0$ such that $e_{\lambda}\varphi(x)>$-$\infty$ for some $x\in{\rm I\!R}^{n}$. The supremum over all such $\lambda$ is the threshold $\lambda_{\varphi}$ of prox-boundedness for $\varphi$. Prox-boundedness admits various equivalent descriptions and ensures a favorable behavior of Moreau envelopes and proximal mappings. We recall the following results from [54, Theorem 1.25 and Example 10.32].
@@
Let $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ be l.s.c. and prox-bounded with threshold $\lambda_{\varphi}>0$. Then for all $\lambda\in(0,\lambda_{\varphi})$, we have:
@@
The next proposition presents some properties of Moreau envelopes and proximal mappings that provides, in particular, the constructive choice of the parameter $\lambda$ of the Moreau envelope (2.12) in comparison with [47, Theorem 4.4], which is important for the design and justification of algorithms.
@@
Let $r\geq 0$ and $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ be prox-bounded, $r$-level prox-regular at $\bar{x}$ for $\bar{v}\in\partial\varphi(\bar{x})$. Then we have
$P:=\left\{\lambda\in\left(0,\frac{1}{r}\right)\Big|\;\varphi(x)\geq\varphi(\bar{x})+\langle\bar{v},x-\bar{x}\rangle-\frac{1}{2\lambda}\|x-\bar{x}\|^{2}\quad\text{for all }\;x\in{\rm I\!R}^{n}\right\}\neq\emptyset$ (2.14)
with the convention that $1/0:=\infty.$ Let $\alpha\in P$ and $\bar{\lambda}\in(0,\alpha)$. Moreover, for all $\lambda\in\left(0,\bar{\lambda}\right)$, there exists a neighborhood $U_{\lambda}$ of $\bar{x}+\lambda\bar{v}$ on which the following hold:
@@
(ii) The proximal mapping $\text{\rm Prox}_{\lambda\varphi}$ from (2.13) is single-valued and Lipschitz continuous on $U_{\lambda}$ while satisfying the condition ${\rm Prox}_{\lambda\varphi}(\bar{x}+\lambda\bar{v})=\bar{x}$.
@@
$\nabla e_{\lambda}\varphi(x)=\left(\lambda I+(\partial\varphi)^{-1}\right)^{-1}(x)\quad\text{for all}\quad x\in U_{\lambda}.$ (2.16)
[Proof.] First we show that $P\neq\emptyset$ for $P$ from (2.14). Indeed, the $r$-level prox-regularity of $\varphi$ at $\bar{x}$ for $\bar{v}$ yields
$\varphi(x)\geq\varphi(\bar{x})+\langle\bar{v},x-\bar{x}\rangle-\frac{r}{2}\|x-\bar{x}\|^{2}\quad\text{for all }\;x\;\text{sufficiently near }\;\bar{x}.$
Combining the latter with the prox-boundedness of $\varphi$, we deduce from the proof of [54, Proposition 8.46(f)] that there exists $\widetilde{\lambda}\in(0,1/r)$ such that
$\varphi(x)\geq\varphi(\bar{x})+\langle\bar{v},x-\bar{x}\rangle-\frac{1}{2\widetilde{\lambda}}\|x-\bar{x}\|^{2}\quad\text{for all }\;x\in{\rm I\!R}^{n},$
which ensures that $P\neq\emptyset$ and thus justifies (2.14). Next we show that assertions (i)–(iii) hold for all $\lambda\in\left(0,\bar{\lambda}\right)$. To this ens, fix such $\lambda$ and consider the function $\phi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ defined by
$\phi(x):=\varphi(x+\bar{x})-\varphi(\bar{x})-\langle\bar{v},x\rangle\quad\text{for all }\;x\in{\rm I\!R}^{n}.$ (2.17)
It is clear that $\phi(0)=0$, $0\in\partial\phi(0)$, and
$\text{\rm Prox}_{\lambda\phi}(x)=\text{\rm Prox}_{\lambda\varphi}(x+\bar{x}+\lambda\bar{v})-\bar{x}\quad\text{and }\;e_{\lambda}\phi(x)=e_{\lambda}\varphi(x+\bar{x}+\lambda\bar{v})-\langle\bar{v},x\rangle-\varphi(\bar{x})-\frac{\lambda}{2}\|\bar{v}\|^{2},$ (2.18)
for all $x\in{\rm I\!R}^{n}.$ Since $\alpha\in P$, and $\bar{\lambda}<\alpha$, we deduce from (2.17) that
$\phi(x)>-\frac{1}{2\bar{\lambda}}\|x\|^{2}\quad\text{for all }\;x\neq 0.$ (2.19)
Moreover, the $r$-level prox-regularity of $\varphi$ at $\bar{x}$ for $\bar{v}$ ensures the $r$-level prox-regularity of $\phi$ at 0 for 0, which also tells us that $\phi$ is $1/{\bar{\lambda}}$-level prox-regular at 0 for 0. Combining this with (2.19) and applying [47, Theorem 4.4] to the function $\phi$, we can find a neighborhood $X_{\lambda}$ of 0 such that $\text{\rm Prox}_{\lambda\phi}$ is single-valued and Lipschitz continuous on $X_{\lambda}$ with $\text{\rm Prox}_{\lambda\phi}(0)=0$, while the Moreau envelope $e_{\lambda}\phi$ is of class $\mathcal{C}^{1,1}$ on $X_{\lambda}$ with
$\nabla e_{\lambda}\phi(x)=\frac{1}{\lambda}\left(x-\text{\rm Prox}_{\lambda\phi}(x)\right)\quad\text{for all }\;x\in X_{\lambda}.$ (2.20)
Moreover, the following equation holds if $\phi$ is subdifferentially continuous at 0 for 0:
$\nabla e_{\lambda}\phi(x)=(\lambda I+(\partial\phi)^{-1})^{-1}(x)\quad\text{for all }\;x\in X_{\lambda}.$ (2.21)
Defining $L_{\lambda}(x):=x+\bar{x}+\lambda\bar{v}$ for all $x\in{\rm I\!R}^{n}$, we deduce that $U_{\lambda}:=L_{\lambda}(X_{\lambda})$ is a neighborhood of $\bar{x}+\lambda\bar{v}$. Fix an arbitrary $x\in U_{\lambda}$ and find $x^{\prime}\in X_{\lambda}$ such that $x=x^{\prime}+\bar{x}+\lambda\bar{v}$. By (2.18), it follows that
$\text{\rm Prox}_{\lambda\varphi}(x)=\text{\rm Prox}_{\lambda\phi}(x^{\prime})+\bar{x}\quad\text{and }\;\nabla e_{\lambda}\varphi(x)=\nabla e_{\lambda}\phi(x^{\prime})+\bar{v}.$
Combining the latter with (2.20), we have
$\nabla e_{\lambda}\varphi(x)=\nabla e_{\lambda}\phi(x^{\prime})+\bar{v}=\frac{1}{\lambda}\big(x^{\prime}-\text{\rm Prox}_{\lambda\phi}(x^{\prime})\big)+\bar{v}=\frac{1}{\lambda}\big(x-\text{\rm Prox}_{\lambda\varphi}(x)\big),$ (2.22)
which verifies (2.15). Supposing further that $\varphi$ is subdifferentially continuous at $\bar{x}$ for $\bar{v}$, we show that (2.16) holds. Indeed, due to (2.17), $\phi$ is also subdifferentially continuous at 0 for 0, and thus we deduce from (2.21) that $\nabla e_{\lambda}\phi(x^{\prime})=(\lambda I+(\partial\phi)^{-1})^{-1}(x^{\prime})$. This implies that
$\nabla e_{\lambda}\varphi(x)=\nabla e_{\lambda}\phi(x^{\prime})+\bar{v}\in\partial\phi(x^{\prime}-\lambda\nabla e_{\lambda}\phi(x^{\prime}))+\bar{v}=\partial\varphi(\bar{x}+x^{\prime}-\lambda\nabla e_{\lambda}\phi(x^{\prime}))=\partial\varphi(x-\lambda\nabla e_{\lambda}\varphi(x)),$
which tells us in turn that
$\nabla e_{\lambda}\varphi(x)\in(\lambda I+(\partial\varphi)^{-1})^{-1}(x)\neq\emptyset.$ (2.23)
Let us next verifies that $(\lambda I+(\partial\varphi)^{-1})^{-1}(x)$ is singleton. Suppose that $v_{1},v_{2}\in(\lambda I+(\partial\varphi)^{-1})^{-1}(x)$ and get that $v_{1}\in\partial\varphi(x-\lambda v_{1})$ and $v_{2}\in\partial\varphi(x-\lambda v_{2})$. The continuous $r$-level prox-regularity of $\varphi$ at $\bar{x}$ for $\bar{v}$ yields
$\langle x-\lambda v_{1}-x+\lambda v_{2},v_{1}-v_{2}\rangle\geq-r\|x-\lambda v_{1}-x+\lambda v_{2}\|^{2},$
or $(1-\lambda r)\|v_{1}-v_{2}\|^{2}\leq 0$, which tells us that $v_{1}=v_{2}$ since $\lambda\in(0,1/r)$. Therefore, $(\lambda I+(\partial\varphi)^{-1})^{-1}(x)$ is singleton. Combining the latter with (2.23), we obtain (2.16) and thus complete the proof.
Remark 2.6 (on the choice of the Moreau envelope parameter).
As in Proposition 2.5, the parameter $\lambda$ ensuring the fulfillment of the assertions in this result depends on the value of $\bar{\lambda}$. The determination of $\lambda$ can be simplified in the following situations:
(i) When $\bar{x}\in\text{\rm argmin}\,\varphi$, we have $\varphi(x)\geq\varphi(\bar{x})$ for all $x\in{\rm I\!R}^{n}$. In this case, it follows that $P=(0,1/r)$ and thus $\lambda$ can be chosen arbitrarily in $(0,1/r)$, where $r$ is the level of prox-regularity of $\varphi$ at $\bar{x}$ for $\bar{v}$.
(ii) When $\varphi$ is $r$-weakly convex, we always have
$\varphi(y)\geq\varphi(x)+\langle v,y-x\rangle-\frac{r}{2}\|y-x\|^{2}\quad\text{for all }x,y\in{\rm I\!R}^{n},$
and hence $\lambda$ can be chosen in $(0,1/r)$. In the special case where $\varphi$ is convex, we have $r=0$, which implies that $\lambda$ can be chosen arbitrarily in $(0,\infty)$.
(iii) When $\bar{x}\in\text{\rm Prox}_{\alpha\varphi}(\bar{x}+\alpha\bar{v})$ for some $\alpha>0$, we have $\alpha\in P$. Indeed, it follows from (2.13) that
$\varphi(x)+\frac{1}{2\alpha}\|x-\bar{x}-\alpha\bar{v}\|^{2}\geq\varphi(\bar{x})+\frac{1}{2\alpha}\|\bar{x}-\bar{x}-\alpha\bar{v}\|^{2}\quad\text{for all }\;x\in{\rm I\!R}^{n},$
which is equivalent to
$\varphi(x)\geq\varphi(\bar{x})+\langle\bar{v},x-\bar{x}\rangle-\frac{1}{2\alpha}\|x-\bar{x}\|^{2}\quad x\in{\rm I\!R}^{n},$
and thus $\alpha\in P$. In this case, we can choose $\lambda\in(0,\alpha)$ ensuring the fulfillment of Proposition 2.5.
@@
Let $\sigma>0$ and $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ be continuously prox-regular at $\bar{x}\in\mbox{\rm dom}\,\varphi$ for $0\in\partial\varphi(\bar{x})$. The following assertions are equivalent:
(i) $\varphi$ is $\sigma$-strongly variationally convex at $\bar{x}$ for $\bar{v}=0$.
@@
Then $\varphi$ is twice epi-differentiable at $\bar{x}$ for $v$ if for every $w\in{\rm I\!R}^{n}$ and $\tau_{k}\downarrow 0$ there is a sequence $w^{k}\to w$ with $\Delta^{2}_{\tau_{k}}\varphi(\bar{x},v)(w^{k})\to d^{2}\varphi(\bar{x},v)(w)\;\mbox{ as }\;k\to\infty.$ Twice epi-differentiability has been recognized as an important concept of second-order variational analysis with numerous applications to optimization; see, e.g., [54]. To formulate the related notion of generalized twice differentiability of functions, we proceed following [51]. Given a quadratic matrix $A\in{\rm I\!R}^{n\times n}$, denote the quadratic form associated with the matrix $A$ by
$q_{A}(w):=\langle w,Aw\rangle\;\mbox{ for all }\;w\in{\rm I\!R}^{n}.$
@@
$\Omega_{\varphi}:=\big\{(x,v)\in\mbox{\rm gph}\,\partial\varphi\;\big|\;\varphi\;\text{ is generalized twice differentiable at }x\;\text{for }v\big\}.$
@@
$\mbox{\rm gph}\,\partial q\subset\mbox{\rm gph}\,\partial^{2}\varphi(\bar{x},\bar{v})\quad\text{for any }\;q\in\mathrm{quad}\,\varphi(\bar{x}|\bar{v}).$